### v3

# Label-efficient Time Series Representation Learning: A Review

Abstract
The scarcity of labeled data poses a significant challenge to deploying deep learning models for time series data in real-world applications. To address this, various strategies, e.g., transfer learning, self-supervised learning, and semi-supervised learning, have been developed to enhance the learning capabilities of these models using limited labeled time series data. In this survey, we introduce a novel taxonomy for the first time, categorizing existing approaches based on their reliance on external data sources to tackle the labeled data scarcity issue in time series. Furthermore, we present a review of the recent advances in each strategy, conclude the limitations of current methodologies, and suggest future research directions that promise further improvements in the field.
Index Terms:
Time series, Label-efficient learning, Data augmentation, Self-supervised learning, Semi-supervised learning, and Domain adaptation.
{impact}
This research demystifies the application of deep learning to time series data with limited labels. It sets the stage for significant technological advancements, enabling improved model performance in healthcare, finance, and environmental monitoring without the extensive need for labeled data. Economically, it offers a pathway to cost savings in data processing, while socially, it could improve public health and environmental management. By simplifying complex AI methodologies for non-specialists, this work not only progresses the AI field but promises widespread benefits across various sectors, making deep learning more accessible and impactful in real-world applications.
Figure 1: Graphical overview of the different scenarios of data availability and the suitable solution. Note that the directions given by arrows are not the only solutions for each respective scenario, but can be the best way to achieve the best performance.
I Introduction
Deep learning has shown significant success in diverse time series applications. Unfortunately, the applicability of deep learning models in real-world scenarios is tied to the availability of large and labeled datasets, which are usually lacking for time series applications [1, 2]. For example, we can collect massive amounts of sensory data from manufacturing sensors, while only a few of them are being annotated by domain experts. Also, in some healthcare applications, it can be difficult to collect or access enough data, due to security reasons or the patient’s right to privacy, where we end up with only a few (labeled) samples [3]. In either situation, training efficient deep learning models with these data and annotation limitations can be very challenging.
Over recent years, a significant body of research has emerged to address the label scarcity challenges, with methodologies spanning various conceptual paradigms [4, 5]. These methodologies can be categorized into in-domain methods and cross-domain methods. In-domain methods focus on deriving representations from the available time series data, without the reliance on external data sources. Strategies within this category include data augmentation, which increases the number of labeled samples. Additionally, self-supervised learning allows learning from fully unlabeled data [6], while semi-supervised learning provides a harmonized approach to learning from both labeled and unlabeled instances within the same domain [7]. Conversely, cross-domain methods learn representations from an external, labeled source domain, subsequently transferring this acquired knowledge to the target domain of interest. The common strategy is to use transfer learning approaches, either by directly fine-tuning the pretrained model on the target domain or by adapting the source and target domains to have similar distributions and class boundaries [5].
In this survey, we navigate through three specific real-world scenarios characterized by the availability of data: scenarios with a limited number of labeled samples, those with exclusively unlabeled time series data, and situations combining both labeled and unlabeled data. We introduce a novel taxonomy that systematically organizes the solutions to these scenarios into in-domain and cross-domain strategies. This taxonomy, depicted in Fig. 1, guides the classification of each solution based on its reliance on external data sources. Our objective extends beyond merely cataloging recent advancements within these categories. We aim to craft a cohesive roadmap that guides the effective deployment of deep learning models despite the constraints posed by limited labeled data. It is important to note that our intention is not to provide an exhaustive discussion of each method within every category. Instead, we strive to offer a panoramic view, presenting a high-level understanding of the challenges and available solutions. By doing so, we enable readers to grasp the overarching concepts and approaches, facilitating further exploration and in-depth investigations according to their specific research interests.
Overall, the contributions of this paper can be summarized as follows:
  • We conduct a comprehensive survey of label-efficient learning techniques tailored for time series data. This includes scenarios with limited labeled samples, entirely unlabeled data, and their combinations, providing a detailed overview of the current landscape and the inherent challenges of limited labeled data.
  • We propose a novel taxonomy that systematically categorizes existing approaches into in-domain and cross-domain strategies, offering a new lens through which to understand and navigate the solutions addressing label scarcity in time series analysis. It underscores the dependency of these strategies on external data sources, presenting a unique framework for future research and application.
  • We evaluate the challenges, emerging trends, and future directions in label-efficient time series analysis.
II Related Surveys
In the realm of label-efficient time series representation learning, several surveys have delved into specific solutions tailored to address the challenges of limited labeled data [8, 9, 10, 4, 11, 5]. While these surveys have significantly contributed to the field, they tend to focus on specific techniques or solutions, limiting their scope to a singular aspect of label-efficient time series representation learning. Our survey, on the other hand, synthesizes these individual perspectives into a cohesive narrative that encompasses a broader range of strategies.
We first discuss surveys related to in-domain solutions. Wen et al. [8] comprehensively summarize time series data augmentation techniques across key tasks such as forecasting, anomaly detection, and classification. The authors systematically review different augmentations from simple time-domain transformations to more intricate methods in transformed frequency and time-frequency domains. Also, Deldari et al. [9] explore the self-supervised representation learning to learn from unlabeled time series data across various modalities, with an emphasis on temporal data and cross-modal learning models. The authors present a classification scheme to benchmark, and categorize various studies, as a guide for model selection.
Furthermore, Zhang et al. [10] present an exhaustive review of state-of-the-art self-supervised learning techniques specifically tailored for time series data. They categorize these methods into three distinct paradigms: generative-based, contrastive-based, and adversarial-based, and they delve into various frameworks and techniques within each category. Similarly, Ma et al. [4] thoroughly examine time series pre-trained models, where they discuss time series modeling tasks and the associated deep learning models. They also discuss supervised, unsupervised, and self-supervised pre-training techniques.
Regarding the cross-domain related surveys, inspired by the success of transfer learning in computer vision, Ismail Fawaz et al. [11] explored the potential of transfer learning for time series classification and conducted exhaustive experiments to draw conclusions. Their findings underscore the presence of transferable low-level features in time series, and their transferability can be measured by the Dynamic Time Warping technique. Last, Shi et al. [5] offer a thorough discussion about unsupervised domain adaptation in the context of time series sensor data. They categorize these studies by industry, detailing application backgrounds, frequently used sensors, data discrepancy sources, and prominent research datasets.
Concluding our discussion on the related surveys, we emphasize that, to our knowledge, our work is the first to comprehensively address label-efficient representation learning for time series data from various perspectives. We provide a broad overview, exploring multiple scenarios and their solutions in depth.
III Preliminaries
III-A Literature Search and Selection Strategy
Our review of time series representation learning for label-efficient scenarios was founded on a systematic search and selection methodology, tailored to capture the recent advances and the relevant contributions to the field.
Search Strategy: We utilized leading academic databases, including IEEE Xplore, PubMed, Google Scholar, and ACM Digital Library, focusing on publications in the last five years. Keywords, e.g., “time series analysis”, “deep learning”, and “semi-supervised learning” guided our search, aiming to include a wide range of studies pertinent to label efficiency in time series data.
Selection Criteria: Articles were chosen based on peer-review status and direct relevance to label-efficient learning in time series analysis. We excluded non-peer-reviewed sources, except for those showing high impact on the field, e.g., [12, 13] or recent surveys. We also excluded unrelated topics, and older publications, unless they offered foundational insights.
Selection Process: An initial screening of titles and abstracts was followed by a full-text review to assess each article’s contribution to the field. This dual-phase process ensured a comprehensive and unbiased selection of literature. This streamlined approach enabled us to distill significant advancements and methodologies in time series representation learning, addressing the challenges of label scarcity.
III-B Representation Learning Problem Formulation
In this section, we present a mathematical framework that encapsulates the essence of time series label-efficient learning, covering both in-domain and cross-domain learning scenarios. This framework will set the stage for a deeper exploration of the specific strategies you mentioned.
Consider a time series dataset $\mathcal{D}$ composed of pairs $(X,Y)$, where $X=\{x_{1},x_{2},...,x_{N}\}$ represents the time series data points and $Y=\{y_{1},y_{2},...,y_{N}\}$ denotes their corresponding labels. In label-efficient learning scenarios, the goal is to maximize the predictive performance of a learning model $f:X\rightarrow Y$ while minimizing the dependence on labeled instances. Given that a time series instance $x_{i}\in X$ is a sequence of observations $x_{i}=\{o_{1},o_{2},...,o_{T}\}$, where $T$ is the length of the series and $o_{t}$ represents the observation at time $t$. For supervised scenarios, each $x_{i}$ has a corresponding label $y_{i}\in Y$. In label-efficient scenarios, $Y$ may be partially observed or observed with noise.
In-Domain Learning: The dataset $\mathcal{D}$ can vary in composition, encompassing labeled $\mathcal{D}_{L}=\{(x_{i},y_{i})\}_{i=1}^{L}$, unlabeled $\mathcal{D}_{U}=\{x_{j}\}_{j=L+1}^{N}$, or a combination of both subsets. The learning objective adapts to this variability, aiming to utilize $\mathcal{D}_{L}$, $\mathcal{D}_{U}$, or their integration to effectively train a model $f$ that can accurately predict $Y$ for new instances, accommodating the dataset’s specific labeling scenario.
Cross-Domain Learning: Here, we consider two datasets: a source domain $\mathcal{D}_{S}=\{(x^{S}_{i},y^{S}_{i})\}$ with abundant labels and a target domain $\mathcal{D}_{T}=\{(x^{T}_{j},y^{T}_{j})\}_{j=1}^{M}$ where labels may be scarce or absent. The objective is to adapt the knowledge from $\mathcal{D}_{S}$ to improve learning on $\mathcal{D}_{T}$.
III-C Deep Learning Architectures
Selecting an appropriate architecture is foundational to the construction of a label-efficient deep learning model. Various deep learning architectures have been adopted, summarized as follows.
III-C1 Convolutional Neural Networks (CNNs)
CNNs, including 1D-CNN [14], 1D-ResNet [15], and temporal convolutional network (TCN) [12], are prevalent in time series classification. Other designs were proposed, introducing different designs, e.g., InceptionTime [16] and Omni-Scale CNN blocks [17]. These networks excel at identifying patterns and trends with fewer parameters, reducing overfitting risks.
III-C2 Recurrent Neural Networks (RNNs)
RNNs, including LSTM models, are tailored for time series data, incorporating memory components to leverage past timesteps [18]. They usually excel in forecasting and can be utilized alongside or following CNNs in classification tasks [19, 20]. While effective in learning from historical data, RNNs struggle with long sequences and are computationally intensive due to their sequential nature.
III-C3 Attention Mechanism
The attention mechanism was proposed to enhance LSTM by focusing on specific input parts and assigning scores to hidden states [21]. Later, it became vital in applications [22, 23, 24]. Recently, the Transformer model became emerging as a prominent architecture for time series [25, 26]. While attention improves long sequence handling and model interpretability, its computational cost and overfitting risk due to more parameters are notable drawbacks.
IV In-domain Representation Learning
In this section, we discuss three approaches to learning representations from the data in hand. First, we assume possessing only a few labeled samples which can be addressed with data augmentation. Second, we assume the existence of a large cohort of unlabeled data and we discuss how self-supervised learning can be utilized. Third, we may have some labeled samples besides the massive unlabeled data, where we can apply the semi-supervised learning techniques.
IV-A Data Augmentation
Data augmentation is usually adopted as a way to artificially scale up the training dataset, by applying several transformations to the input data. This can be particularly useful when it is difficult or expensive to collect more real-world data. Fig. 2 illustrates the principle of data augmentation applied to a single time series example. The different augmentations shown in the figure elucidate how one original sequence can be transformed into multiple variants, each serving as a distinct training example.
Given a labeled dataset $\mathcal{D}_{L}=\{(x_{i},y_{i})\}_{i=1}^{L}$, data augmentation aims to generate an augmented dataset $\mathcal{D}_{\text{aug}}=\{(\tilde{x}_{i},\tilde{y}_{i})\}$ where $\tilde{x}_{i}$ is a modified version of $x_{i}$ obtained through various augmentation techniques (e.g., warping, noise injection), and $\tilde{y}_{i}=y_{i}$ assuming the label is invariant to the augmentation. The objective is to train a model $f$ to improve generalization, and minimize $\mathcal{L}(f(\tilde{x}_{i}),\tilde{y}_{i})$ over $\mathcal{D}_{\text{aug}}\cup\mathcal{D}_{L}$.
By adding more samples, data augmentation aims to provide the model with diverse instances that are representative of the underlying data distribution. These transformed samples not only contribute to better generalization but also alleviate the need for an excessive volume of labeled data. Aside from increasing the sample size, the variety of augmentations applied to the time series data can generate new sequences that are realistic variations of the original data, mimicking natural fluctuations and nuances. Thus, the model can become robust against different orientations, scales, noise, and other variations that might be present in real-world data. In addition, some works utilize data augmentation to upsample the minority classes to overcome the class imbalance issue in some time series applications [27].
Figure 2: Increasing the number of samples through augmentations. For each input sample, we apply three (or more) augmentations to generate new samples
There are several data augmentation techniques proposed for time series data. Wen et al. [8] provide a comprehensive set of time series data augmentations that can be applied in the time domain and the frequency domain. They categorized the augmentations into basic and advanced approaches and provide several subcategories under each, e.g., decomposition methods and statistical generative models. Similarly, Iwana and Uchida [28] conduct a survey of data augmentation for empirically evaluating the classification tasks, where they explore data transformations and the different ways to mix time series patterns. In general, some augmentations are more widely employed, e.g., scaling, shifting, rotation, masking, and noise injection.
Besides applying transformations, synthetic data generation is yet another technique to artificially generate data using simulation. This technique can be particularly useful in preserving data privacy, as it might not be viable to access real-world data in some cases. Therefore, generating synthetic data provides a way to use the data without compromising privacy. There are several different methods to generate synthetic data, but we mainly explore the Generative Adversarial Networks (GANs), being the most commonly used approach. For example, ActivityGAN generates time series sensory data for human activity recognition with 1D-CNN networks by training a GAN [29]. They find that performance increases by adding more synthetic data to real data. Similarly, DCGAN generates artificial EEG signals to improve the performance of the Brain-Computer Interface (BCI) tasks [30], where the generator receives the subject-specific feature vector to generate artificial samples. By training it against the discriminator, it learns to generate artificial samples similar to the original ones. Their model witnessed a significant improvement by including the generated data.
IV-B Self-supervised Learning
Self-supervised learning involves training a model to learn representations from data using only the unlabeled samples. Instead of relying on human-provided labels, we define a new task to be solved by the model or generate pseudo labels based on the input data. The general process of self-supervised learning is shown in Fig. 3. The training process is usually performed in two phases. In the first phase, i.e., the pre-training phase, we feed the encoder model with the unlabeled data and train it with the self-supervised task. Next, we use the few labeled samples in the second phase, i.e., the fine-tuning, to tailor the model to the downstream task, and improve its performance on these few labels, compared to fully supervised training with the few labels.
Self-supervised learning methods can be categorized into auxiliary- or contrastive-based methods. In both categories, the unlabeled data is used to learn an additional task that is not the main focus of the model but can provide additional supervision for learning useful representations of the data.
Figure 3: Pipeline of the self-supervised learning process. The pretraining phase receives the input time series, sends it to the encoder to extract the features, and performs the self-supervised learning task. Next, the pretrained encoder is fine-tuned with the available few labeled data.
IV-B1 Auxiliary-based Methods
The auxiliary-based methods propose an auxiliary task that can be related to the downstream task or the characteristics of the data, which helps the pretrained model to perform well on the downstream task. Given an unlabeled dataset $\mathcal{D}_{U}=\{x_{j}\}_{j=L+1}^{N}$, we define a pretext task that generates pseudo-labels $\hat{y}_{j}$ for $x_{j}$. The model $f$ learns to predict $\hat{y}_{j}$ from $x_{j}$, effectively learning useful representations without manual labeling.
For example, Banville et al. [31] introduce two auxiliary tasks, i.e., relative positioning and temporal shuffling, to learn the underlying temporal characteristics in the EEG signals. In the first task, a window of the signal is used as an anchor to measure the relative position of other windows, while in the second task, a window of the signal is shuffled to teach the model to fix the temporal order of the signal. Similarly, Saeed et al. [32] propose some auxiliary tasks, e.g., feature prediction from a masked window, temporal shift prediction, modality denoising, and blend detection, to learn representations from human activity recognition data. In addition, TST [25] and STraTS [33] propose to predict the masked part in the signal as an auxiliary task to pre-train a Transformer model. Furthermore, Islam and Washington [34] introduce a personalized stress prediction system that utilizes wearable multimodal biosignal time series data. Their approach relies on a self-supervised forecasting task for pre-training models on individual subjects, then fine-tuning specifically for the task of stress prediction.
Another example of auxiliary tasks is to utilize data augmentations to provide several transformations to the time series data. Each transformation is assigned a pseudo label, and the model is trained to classify them with a standard cross-entropy loss. Several works follow this approach in different applications, e.g., electrocardiogram (ECG) emotion recognition [35] and human activity recognition [36].
Even though auxiliary tasks have shown effectiveness, they restrict the generality of the learned representations [13]. Therefore, most time series self-supervised learning works have recently adopted contrastive learning instead [37].
IV-B2 Contrastive-based Methods
Contrastive-based self-supervised learning methods aim to form positive and negative pairs, where the contrastive loss maximizes the similarity among positive pairs while minimizing the similarity among negative pairs. To illustrate, given an unlabeled dataset $\mathcal{D}_{U}=\{x_{j}\}_{j=1}^{N}$ of time series instances, the objective of contrastive self-supervised learning is to learn a representation $f$ that embeds similar instances, i.e., positive pairs closer together and dissimilar instances, i.e., negative pairs, further apart in the embedding space. The positive pairs are data instances $(x_{i},x_{i}^{+})$ derived from the same time series instance or through augmentations that preserve the semantic content (see Section IV-A). The negative pairs are $(x_{i},x_{j}^{-})$, where $x_{i}$ and $x_{j}^{-}$ do not share the same semantic content, often chosen from different time series instances within the dataset. The objective is to learn a function $f$ that minimizes a contrastive loss, such as the InfoNCE loss, defined for a batch of instances as follows:
$\mathcal{L}_{\text{contrastive}}=-\sum_{i=1}^{n}\log\frac{\exp(sim(f(x_{i}),f(x_{i}^{+}))/\tau)}{\sum_{x_{j}^{-}\in N}\exp(sim(f(x_{i}),f(x_{j}^{-}))/\tau)},$
where $sim(u,v)$ is a similarity measure between two embeddings (e.g., the dot product or cosine similarity), $\tau$ is a temperature parameter that scales the similarity measure. Notably, the sum in the denominator runs over a set of negative pairs for each positive pair, and $n$ is the number of positive pairs in a batch.
Previous works propose different strategies to choose the positive and negative pairs to improve the quality of the learned representations. Additionally, most works attempt to learn the temporal relations within time series in the pretraining phase. In this section, we provide a new time series-specific categorization for contrastive-based self-supervised learning methods. Particularly, we categorize them into intra-sample contrastive methods, inter-sample contrastive methods, and frequency domain contrastive methods.
Inter-sample Contrasting (Contextual Contrasting)
In inter-sample contrastive learning, the model uses the input signal as a whole without splitting or windowing. One of the well-performing visual contrastive methods is SimCLR [6], which relies on data augmentations to define positive and negative pairs. In specific, the positive pairs are the augmented views of an anchor sample, while the negative pairs are all the augmented views of other samples. Many works followed SimCLR for time series data, such as [38, 39] that apply SimCLR-like methods for sleep stage classification and fault diagnosis tasks respectively. In addition, TimeCLR [40] adapts the SimCLR framework besides introducing DTW-based data augmentation and utilizing InceptionTime for feature extraction to tackle the challenge of sparse labels.
ExpCLR [41] leverages expert knowledge instead of relying on data transformations for contrastive learning. Recognizing the availability of expert features in fields like industry and medicine, this method incorporates such features to ensure the development of useful time-series representations that fulfill two proposed essential properties often overlooked by current methods. Mixup Contrastive Learning (MCL) proposes to replace the traditional contrastive loss with a mixup-based contrastive loss [42]. Specifically, MCL generates new mixed samples in each mini-batch using the mixup operation, and the proposed loss predicts the mixup ratio among each pair of samples. Also, Chang et al. [43] propose another loss that searches for positive and negative pairs for each anchor sample in both the input and embedding spaces. The contrastive loss is then applied to the selected triplet to learn a causal temporal CNN.
MHCCL [44] learns semantic-rich representations from unlabeled multivariate time series data, utilizing a Masked Hierarchical Cluster-wise Contrastive Learning framework. It addresses the issue of false negative pairs in traditional contrastive learning by exploiting hierarchical clustering for semantic information at multiple granularities. With downward and upward masking strategies, MHCCL efficiently filters out fake negatives and refines cluster prototypes, enhancing the clustering process and quality. The idea of supervised contrastive loss was also adopted by MICOS [45], where they blend self-supervised, intra-class, and inter-class supervised contrastive learning to effectively utilize available labels to address the challenges in capturing complex spatio-temporal features inherent in multivariate time series. Last, InfoTS proposes a contrastive learning framework that utilizes information theory to guide the selection of data augmentations, ensuring both high fidelity to the original data and sufficient variety to foster robust and discriminative representation learning [46]. By applying information theory, InfoTS quantitatively evaluates potential augmentations based on their ability to preserve crucial information content of the time series data, thereby ensuring the augmented samples are representative and beneficial for learning discriminative features.
Intra-sample Contrasting (Temporal Contrasting)
Intra-sample contrastive methods aim to learn the temporal relations in time series data in the pretraining phase. To do so, they process the signal at the timestep level, either in the embedding space or in the input space. One of the earliest works is Contrastive Predictive Coding (CPC), which learns the temporal relations by splitting the embedding space of the signal into a past part and a future part [13]. Then, an autoregressive model is trained to produce a context vector from the past part and use it to predict the future part using contrastive learning. Similarly, SleepDPC generates multiple context vectors and uses each one to predict the next timestep in the future part for the sleep stage classification task [47]. Also, TS-TCC applies strong and weak augmentations for each input signal, generates the embeddings for each, and uses the past part of each view to predict the future part of the other view in a tough cross-view prediction task [48]. Last, ContrNP replaces the autoregressive model used by CPC with a neural process for time series forecasting [49], effectively bypassing the need for manually designed data augmentations. By leveraging a variety of sampling functions, this approach generates augmented data variants, extending neural processes with a novel contrastive loss to learn time series representations in a self-supervised manner.
Other methods learn the temporal relations among timesteps in the input space. For example, Temporal Neighborhood Coding (TNC) assumes that the distribution of one window of timesteps is similar to its neighboring windows and can be more distinguishable than the non-neighboring windows [50]. Therefore, TNC forms positive pairs among neighboring windows and negative pairs among non-neighboring windows. Another example is TS2Vec, which applies data augmentation on the input signal and considers the current timestep from the augmented views as positives, while the different timesteps from the same time series are considered as negatives [51]. TS2Vec utilizes a hierarchical contrastive learning approach across augmented context views for robust contextual timestamp representations, which allows for simple aggregation to represent any sub-sequence within a time series.
Frequency Domain Contrastive Methods
A recent trend in time series contrastive learning is to utilize the frequency domain characteristics in representation learning, being one main distinguishing properties of time series data. For instance, Bilinear Temporal-Spectral Fusion (BTSF) applies an iterative bilinear fusion between feature embeddings of both time and frequency representations of time series [52]. Specifically, BTSF generates different views of time series and applies iterative aggregation modules to perform cyclic refinement to the temporal and spectral features. Similarly, both STF-CSL [53] and TF-C [54] apply time domain and frequency domain augmentations, and push the time domain and frequency domain representations of the same sample closer to each other while pushing them apart from representations of other signals. In particular, STF-CSL [53] integrates Short-Time Fourier Neural Networks (STFNets) as the core component and applies both time-domain and frequency-domain data augmentations. STF-CSL aligns with the physical reality of IoT sensing scenarios while also offering substantial performance improvements over traditional time-domain-based self-supervised methods in tasks such as human activity recognition. TF-C [54] introduces a pre-training model focused on overcoming the challenges associated with mismatches in temporal dynamics between pre-training and target domains in time series data. TF-C embeds time-based and frequency-based representations of an example closely together in the time-frequency space, without requiring access to target domain examples during pre-training.
IV-C Semi-supervised Learning
Semi-supervised learning is a type of learning that involves training a model using both labeled and unlabeled data. Combining a small labeled dataset $\mathcal{D}_{L}$ with a larger unlabeled dataset $\mathcal{D}_{U}$, the goal is to train a model $f$ that leverages the unlabeled data $\mathcal{D}_{U}$ alongside $\mathcal{D}_{L}$ to enhance learning. The learning objective often combines a supervised loss $\mathcal{L}_{\text{sup}}(f(x_{i}),y_{i})$ on $\mathcal{D}_{L}$ with an unsupervised loss $\mathcal{L}_{\text{unsup}}(f(x_{j}))$ on $\mathcal{D}_{U}$, where $\mathcal{L}_{\text{unsup}}$ encourages the model to learn meaningful structures from $\mathcal{D}_{U}$.
Semi-supervised learning is particularly useful in situations where it is difficult or expensive to obtain a large amount of labeled data, and only a few labeled samples are available with a large amount of unlabeled data. One of the challenges in semi-supervised learning is how to achieve the best balance of learning from both labeled and unlabeled samples. Fig. 4 provides a visual illustration of the semi-supervised training process. In specific, we feed the model with both the labeled and unlabeled samples, apply the traditional cross-entropy classification loss to the labeled data, and train the model on the unlabeled data with an unsupervised learning technique. In this survey, we categorize semi-supervised learning approaches into self-training-based methods and generic regularization-based methods.
IV-C1 Self-training Methods
Self-training involves assigning labels to the unlabeled data to train the model. It can be implemented either by pseudo-labeling or self-supervised pretraining.
Pseudo-Labeling
[…]
For example, Wang et al. [55] generate pseudo labels from the model predictions to the unlabeled data and use least-square regression to learn both shapelets and classification boundaries. Another way is to use clustering techniques as in SUCCESS, which proposes a hierarchical clustering scheme to assign pseudo labels to samples according to their respective clusters [56]. It calculates the distance to centroids based on the dynamic time warping (DTW) distance metric.
A different way of utilizing the pseudo labels is proposed in FixMatch, which applies two augmentations to the image data, where one augmentation is weak and the other is strong [7]. Then, if the prediction of the weak-augmented view is beyond a threshold value, it uses its pseudo label for the strong augmented view. By showing success with image data, several methods applied FixMatch-like approaches on time series data, e.g., PARSE [57] for EEG-based emotion recognition. PARSE employs pairwise representation alignment to bridge the gap between large volumes of unlabeled data and limited labeled data through a sequence of data augmentation, label guessing and sharpening, and representation alignment. Also, SelfMatch uses signals with Gaussian noise as a weak augmentation and timecut as a strong augmentation [58]. It also applies self-distillation to guide the lower-level blocks of the feature extractor with the knowledge obtained in the output layer.
Figure 4: Illustration of semi-supervised learning process. Usually, the classification cross-entropy loss is applied to the labeled portion of data, while an unsupervised learning technique is applied to learn useful and supportive representations from the unlabeled data.
TABLE I: Summary of In-domain Methods for Time Series Representation Learning.
c—c—c—p1.7cm—p4.9cm—p4cm—p4cm Category   Method Description Advantages Potential Limitations
Data Augmentation [8], [28] Applies transformations, e.g., scaling and masking to generate realistic variations of original time series data Increases model robustness, alleviates need for labeled data May introduce artificial bias if not carefully managed
ActivityGAN [29] Uses GAN to generate sensory data for human activity recognition Helps in data privacy, improves performance with synthetic data Quality of generated data heavily depends on GAN’s training
DCGAN [30] Generates artificial EEG signals to aid BCI tasks Improves performance with the inclusion of generated data Risk of generating non-representative or biased data
Self-supervised Learning Auxiliary-based Methods [31] Introduces tasks to understand EEG signal’s temporal dynamics Enhances temporal understanding May limit generality
Sense and learn [32] Employs various tasks for representation learning in HAR Diverse methods to handle data Potentially restricts representation generality
TST [25], STraTS [33] Pre-trains Transformers by predicting masked signal parts Benefits from parallel processing of Transformer architecture Focused on specific model structures
[35],[36] Trains models to recognize data augmentations Exploits augmentation diversity to find best performing ones Depends on augmentation relevance
Contrastive-based Methods Inter-sample [6], [38], [39] Applies SimCLR to time series Enhances inter-sample learning Requires careful pair selection
MCL [42] Uses mixup operation for contrastive loss Novel mixup application Mixup parameters are critical
[43] Searches for positive and negative pairs in input and embedding spaces Promotes causal learning Complex pair selection process
Intra-sample CPC [13] Learns temporal relations by predicting future from past parts in embedding space Facilitates understanding of temporal dynamics May neglect non-temporal features
SleepDPC [47] Uses multiple context vectors for future timestep prediction in sleep staging Enhances prediction accuracy for specific tasks Task-specific design limits general applicability
TS-TCC [48] Applies cross-view prediction tasks with strong and weak augmentations Improves robustness to augmentation variability Requires careful design for data augmentations
ExpCLR [41] Leverages expert knowledge for contrastive learning focusing on overlooked properties Utilizes domain-specific insights effectively May require access to robust expert features
MHCCL [44] Employs Masked Hierarchical Cluster-wise Contrastive Learning addressing the issue of false negative pairs Enhances representation quality by refining cluster prototypes Complexity in implementing hierarchical clustering and masking strategies
MICOS [45] Combines intra- and inter-class SSL for capturing spatio-temporal features Effectively utilizes available labels for improved learning Challenge to balance the different components of the loss function
ContrNP [49] Replaces autoregressive model with a neural process for prediction Offers flexible context-based predictions Complexity in model training and tuning
Frequency BTSF [52] Utilizes iterative bilinear fusion for time and frequency domain features Enhances feature representation through fusion May be computationally intensive
STFNets [53] Combines time and frequency domain augmentations for representation learning Balances temporal and spectral feature learning Requires careful balance of domain influences
TF-C [54] Enforces consistency between time and frequency domain representations Promotes comprehensive feature representation Complexity in achieving domain consistency
Semi-supervised Learning Self-training Pseudo-labeling [55], [56] Utilizes model predictions to generate pseudo labels for unlabeled data Exploits unlabeled data to improve learning Risk of reinforcing incorrect labels
[7], [57], [58] Applies weak and strong augmentations to generate pseudo labels, leveraging self-distillation Enhances label efficiency and model robustness Dependent on thresholding and augmentation quality
Self-supervised [59], [60] Leverages forecasting and fine-tuning on labeled data for semi-supervised learning Provides robust feature representations May require extensive pretraining
[61], [62], [63], [64] Uses contrastive learning and transformation prediction for semi-supervised learning from time series data Improves learning from unlabeled data Complexity in implementation and tuning
Generic Regularization Ensemble [65], [66] Employs temporal ensembling and multi-modal classifiers for leveraging unlabeled data in learning Reduces the dependency on large labeled datasets May introduce ensemble disagreement
Reconstruction HCAE [67], AnoVAE [68], SMATE [69] Trains models to reconstruct input data, leveraging autoencoders and variational autoencoders for semi-supervised learning Encourages learning of useful representations Reconstruction quality directly impacts performance
REG-GAN [70], SRGAN [71] Utilizes GANs for signal regeneration and learning, applying supervised regression on labeled parts Facilitates learning from unlabeled data GAN training stability issues